OpenStudy (anonymous):

PLEASE SHOW WORK AND HELP!! 3^√45 * 3^√75

OpenStudy (anonymous):

3 ^ (6.70820393) * 3 ^ (8.66025404)

OpenStudy (anonymous):

3 ^ 6.70820393 = 1587.1818 3 ^8.66025404 = 13551.6218

OpenStudy (anonymous):

1587.1818 * 13551.6218 = 21508887.5

OpenStudy (anonymous):

335√+53√

OpenStudy (anonymous):

wth? did u multiply & simplify cuz my final answer was 15

OpenStudy (anonymous):

3^3 5√+53√

OpenStudy (anonymous):

Lol wait, I may have messed up xD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sooo?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

It's weird cause If I put it in wolfram I get the same answer as mine...

OpenStudy (anonymous):

OH wait, It's 2.15?

OpenStudy (anonymous):

nope, I was right. Is that any of your answers?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

wen i put in mathway thats wat is gives me

OpenStudy (anonymous):

Perhaps the problem was entered in error. I would consult your teacher.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

;-; No medal? *sniffle*